Circular No. 8004

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hf IN MCG –05-3-20 Further to IAUC 8002, D. Hutchings and W. D. Li report the LOTOSS

discovery of an apparent supernova (mag ∼ 16.5) on unfiltered KAIT images taken on Oct. 29.3 and 30.3 UT. SN 2002hf is located at α = 0^{h}57^{m}47^{s}.74, δ = −27^{o}30′21′′.5 (equinox 2000.0), which is 10′′.5 east and 15′′.7 south of the nucleus of MCG –05-3-20. KAIT images taken on Oct. 2.3 (limiting mag ∼ 18.5) and 22.3 (limiting mag ∼ 17) showed nothing at this position.

SUPERNOVA 2002hg IN NGC 3306 T. Boles, Coddenham, England, reports the discovery of an apparent

supernova (mag 17.0) on an unfiltered CCD image taken on Oct. 28.219 UT with a 0.35-m reflector in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 10^{h}37^{m}10^{s}.60, δ = +12^{o}39′13′′.2, which is ≈ 6′′.3 east and 3′′.8 north of the center of NGC 3306. M. Schwartz, Cottage Grove, OR, confirmed the new object with the 0.82-m Tenagra telescope on Oct. 30.54 in morning twilight. Boles measured R = 16.7 and position end figures 10^{s}.60 and 12′′.8 from Schwartz’ image. SN 2002hg is not present on Boles’ images from Jan. 16 (limiting mag 18.5), Mar. 2, or Apr. 11, and it is not present on Palomar Sky Survey red (1993) and blue (1994) plates or on a Quick-V survey plate (1985).

SUPERNOVA 2002he IN UGC 4322 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002he (cf. IAUC 8002), obtained by I. Ginsburg on Oct. 30.47 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a possibly peculiar type-Ia supernova near maximum. Some titanium features appear stronger than normal, indicating that this may be a subluminous event. The spectrum is similar to that of SN 2000dk at 4 days before maximum. Adopting the NED recession velocity of 7364 km/s for the host galaxy, the supernova expansion velocity is ∼ 13900 km/s for Si II (rest 635.5 nm).

V4744 SAGITTARII T. Kato, Kyoto University, reports that nothing was visible at the

location of V4744 Sgr on T-Max 400 exposures taken by H. Nishimura, Kakegawa, Shizuoka, Japan, on Oct. 25.40 UT (limiting mag 10.5).

2002 October 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green